I am not sure what equation to use for the following problem:
total spent $3000.00. total items purchased 1,000. Head lamps cost $2.50 each and Tail lamps cost $3.50. How many of each product was purchased?
I am trying this equation but it does not look accurate:
($2.50x+$3.50y)/1000 = $3000.00

Answer by rfer(16322)

h+t=1000, h=1000-t
2.5h+3.5t=3000
2.5(1000-t)+3.5t=3000
2500-2.5t+3.5t=3000
t=500
h=500
500*2.5=$1250
500*3.5=$1750
total $3000